OpenStudy (anonymous):

What is the sum of an 8–term geometric series if the first term is 12 and the last term is –3,359,232

OpenStudy (mertsj):

What is the formula for the sum of a geometric sequence?

OpenStudy (roadjester):

Plus, you need the common ratio

OpenStudy (anonymous):

I dont know, if I knew that I wouldnt be asking, like if I has a formula I could answer it myself

OpenStudy (mertsj):

Perhaps you good google it. It isn't hard to find.

OpenStudy (roadjester):

great, well we know it's alternating

OpenStudy (anonymous):

\[u_1=12\]\[u_8=-3359232\]\[u_n=u_1*r^{n-1}\]\[S_n=\frac{u_1(1-r^n)}{1-r}\]Here are the formulas. What you want to find is S_8.

OpenStudy (anonymous):

If there's anything you don't get, ask :)

OpenStudy (anonymous):

okay thank you!(: I will

OpenStudy (roadjester):

I've got one. What does u stand for?

OpenStudy (anonymous):

Whats r?

OpenStudy (roadjester):

common ratio

OpenStudy (anonymous):

u is the term. so u_1 is the first term, which is in this case 12. and u_8 is the 8th term. r is the common ratio

OpenStudy (anonymous):

how do I get r?

OpenStudy (anonymous):

from the formula I gave you u_n = u1*r^(n-1)

OpenStudy (roadjester):

r=ratio you get the common ratio by \[\frac {a_{n+1}} a_n\]

OpenStudy (roadjester):

oh, Hairt used u, I use a

OpenStudy (anonymous):

So, the 8th term is -3359232 meaning u_8=-3358232. But u_8 is also: u_8 = u1*r^7= -3359232

OpenStudy (anonymous):

You know u_1 so you can find r

OpenStudy (anonymous):

I actually think its -3359230

OpenStudy (roadjester):

But that's only true for a geometric series.

OpenStudy (anonymous):

okay so can you tell me what formula I use for the problem What is the sum of an 8–term geometric series if the first term is –11, the last term is 180,224, and the common ratio is –4

OpenStudy (anonymous):

Yeah, and this IS a geometric sequence/series

OpenStudy (roadjester):

plug in the numbers...

OpenStudy (roadjester):

a geometric sequence is \[ NOT\] a geometric series

OpenStudy (anonymous):

into what formula?? really, I not trying to be rude, but my head hurts and I am just asking for a formula not answers so please, dont tell me what I ALREADy know

OpenStudy (roadjester):

The formula is at the top, @HairtUB gave you the formula. u1=-11, r=4, and since it's eight terms, n=8

OpenStudy (anonymous):

someone could have said you use the same formula, thanks

OpenStudy (roadjester):

I thought it was self-explanatory...

OpenStudy (roadjester):

my bad

OpenStudy (anonymous):

exactly...